# Inhomogeneity, Fluctuations, and Gap Filling in Overdoped Cuprates

### v1

@@ Abstract
Discussion. As mentioned above, we attribute the smaller subgap to proximity coupling between the large-gap region where the spectrum is measured, and neighboring regions with smaller gaps (see also [46]). Two-gap structures like those obtained here have been observed repeatedly at low temperatures in Pb- and La-doped Bi-2201 [3, 4, 5, 6] and Bi-2212 [7, 47, 8]. In those experiments, the subgap was interpreted as a superconducting gap and the larger gap was attributed to a non-superconducting pseudogap. However, the STM data shows that there are small-gap regions whose coherence peaks coincide with the shoulders observed in large-gap regions [6], which is a key requirement for proximity-induced subgaps. Our calculations predict that these subgaps will fill in when thermal fluctuations wipe out the superconductivity in small-gap regions; that is, the same proximity physics that induces subgaps at low $T$ also wipes them out at high $T$. Conversely, the high-energy coherence peaks, that have been taken for “pseudogaps”, should remain above $T_{c}$, reflecting a robust and phase-coherent local superconducting patch. Unfortunately, detailed measurements of the $T$-dependent STM spectra binned according to gap size have not been reported for the overdoped materials.
@@
We remark that two-gap structures are also ubiquitous in underdoped Bi-2212 [48, 8] and Bi-2201 [4, 11]. Superficially, the measured STS spectra look very much like the results presented here. However, these are complicated by the presence of a pseudogap, charge order, and possibly pair density waves [49]. In particular, it is likely that short-range charge- [50, 51] or pair-density [52, 53] order contribute to the LDOS. However, the mechanism is not understood and simple models of intertwined order generally do not reproduce the measured two-gap spectrum [51] (although see [52, 53]), and we suggest this points to the importance of pairing inhomogeneity in underdoped cuprates. Nonsuperconducting order disappears rapidly with overdoping Bi-2201 [5] and Bi-2212 [54, 55], where our model applies without modification.
@@
One of the remarkable implications of this work is that the pairing interaction is both local and varies by an order of magnitude over nm length scales. This can occur, for example, if the pairing bosons are spin fluctuations and the system is near a spin-density wave transition; in that case, inhomogeneity can create large local differences in the dynamical spin susceptibility [56].
@@
It was recently shown that a spatially random interaction can produce a “strange-metal” linear-in-$T$ resistivity [57, 58, 59] like that measured in optimally and overdoped cuprates. Our work suggests that this same strange-metal physics is responsible for anomalous features in the superconducting spectrum.
@@
To the extent that our model describes them, overdoped cuprates should exhibit aspects of both BKT and granular superconductivity. The superconducting transition in our model is essentially BKT-like, but with an inhomogeneous distribution of local phase-disordering temperatures[29] and gaps[60] that broadens the transition. Our use of the Nunner model [43] for pair inhomogeneity leads to a dilute concentration of nanoscale regions with high mean-field transition temperatures. (The importance of pair-breaking disorder for the transition in overdoped cuprates is debated [60, 1, 2, 23, 25, 61].) In layered materials, the BKT transition is further broadened by interlayer coupling [31].
@@
Thus, while superconductivity disappears via a smeared BKT transition, a weak diamagnetic signal will persist well above $T_{c}$ due to these nanoscale regions. This provides a natural resolution to an apparent dichotomy between magnetoresistance [17] and magnetic susceptibility [18, 20] measurements. The magnetoresistance measurements, which probe the entire sample, will depend on diamagnetic fluctuations in typical superconducting regions, while the onset temperature for a diamagnetic response in susceptibility measurements will depend on the GL temperature of the large-gap regions. We therefore believe that the anomalous spectral properties of the overdoped phase, as measured on inhomogenous families of cuprates, can be explained by our analysis.
I Acknowledgments
@@
M.A.S. is grateful to the late J. Zaanen for numerous inspiring discussions and collaborations on related projects. W.A.A. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). This work was made possible by the facilities of the Shared Hierarchical Academic Research Computing Network (www.sharcnet.ca) and the Digital Research Alliance of Canada. P.J.H. was supported by NSF-DMR-2231821. A.K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT.
Appendix A Supplemental Information
A.1 Bogoliubov-de Gennes Calculations
We model the overdoped cuprate by a square-lattice tight-binding Hamiltonian with $d$-wave pairing treated within mean-field theory:
[…]
where $t_{ij}=t\delta_{\langle i,j\rangle}+t^{\prime}\delta_{\langle\langle i,j\rangle\rangle}-\mu\delta_{i,j}$ with $\delta_{\langle i,j\rangle}$ and $\delta_{\langle\langle i,j\rangle\rangle}$ equal to one for nearest- and next-nearest neighbors, respectively, and zero otherwise. The nearest-neighbor hopping, $t=-1$, sets the energy scale, $t^{\prime}=0.35$, and $\mu$ is adjusted to maintain an electron density $n=0.85$ per unit cell. Note that the value of $n$ does not matter for this work so long as the Fermi surface is qualitatively reasonable. The order parameter $\Delta_{ij}$ is determined self-consistently from
$\Delta_{ij}=V_{ij}\langle c_{i\uparrow}c_{j\downarrow}\rangle,$ (3)
where $i$ and $j$ are nearest-neighbor sites. Self-consistent calculations are performed for $L\times L$ lattices, with $L=40$ with periodic boundary conditions.
@@ A.1 Bogoliubov-de Gennes Calculations
The pairing interaction is assumed to be inhomogeneous because of local modulations due to out-of-plane dopant atoms [43]. We define, at each lattice site,
$V_{i}=V_{0}+\sum^{N_{i}}_{j=1}\pm V_{I}\frac{e^{-\frac{\sqrt{|\mathbf{r}_{i}-\mathbf{r}_{j}|^{2}+r^{2}_{z}}}{\lambda}}}{\sqrt{|\mathbf{r}_{i}-\mathbf{r}_{j}|^{2}+r^{2}_{z}}}.$ (4)
@@
where $N_{i}$ is the number of impurities, the sign ($\pm$) is selected randomly for each impurity, $V_{0}=0.8$, $V_{I}=3.125$, $r_{z}=0.7$, and $\lambda=0.7$. For these calculations, we take $N_{i}=0.075L^{2}$.
@@
where we sum over $N_{k}=15\times 15$ supercells to minimize finite-size effects [23]. In Eq. (5), the $\delta$-function is a Lorentzenian of width $\eta=0.005$.
A.2 Time-Dependent Landau-Ginzburg Calculations
@@
where ${\cal F}$ is the GL free energy, $\Gamma$ is a relaxation rate, and $\zeta_{i}$ is a site-dependent gaussian-distributed complex random variable, taken from the distribution
@@
In this expression, $a_{0}$, $d$, $\Delta_{0}$, and $\xi_{d}$ are GL parameters, $T_{0}$ is a characteristic temperature scale for pairing, and $T_{i\alpha}^{\ast}$ is the local GL temperature below which the singlet pairing interaction along the bond $(i,i+\alpha)$ becomes attractive. Note, however, that $T_{i\alpha}^{\ast}$ is not the local critical temperature, even at the mean-field level, because the order parameter on the bond $(i,i+\alpha)$ depends on the pairing interaction in the neighborhood of the bond. The parameter $d$ determines whether the dominant pairing channel has $d$-wave ($d>0$) or extended $s$-wave ($d<0$) symmetry. In the former case, the lowest-energy GL solution for a homogeneous pairing interaction ($T_{i\alpha}^{\ast}=T_{0}$) has $d$-wave symmetry with $\eta_{x}=-\eta_{y}=\Delta_{0}$ at $T=0$. An extended $s$-wave solution is also possible, with $\eta_{x}=\eta_{y}=\Delta_{0}\sqrt{(a_{0}-2d)/a_{0}}$, however this has higher energy. Values for the GL parameters are shown in Table 1 Both the temperature scale $T_{0}$ and the pairing scale $\Delta_{0}$ are somewhat large for a quantitative description of overdoped cuprates; however, these choices help mitigate finite size effects in our simulations.
Parameter Value Parameter Value
$a_{0}$ 3 $\mu$1.1
$T_{0}$ 0.01 $t$1
$\Delta_{0}$ 0.04 $t^{\prime}$ 0.4
$d$ 3 $t^{\prime\prime}$0.05
$\Gamma dt$ $10^{-3}$
Table 1: Model Parameters. All parameters are written in units of the nearest-neighbour hopping $t$, which in cuprates is $\sim 150$ meV. These parameters correspond to a filling $n=0.85$ and a low-$T$ correlation length $\xi$ of approximately 3 unit cells.
@@
In Eq. (12), the hopping matrix elements $t_{ij}$ are those for a typical cuprate, namely
$t_{ij}=-\mu\delta_{i,j}+t\delta_{\langle i,j\rangle}+t^{\prime}\delta_{\langle\langle i,j\rangle\rangle}+t^{\prime\prime}\delta_{\langle\langle\langle i,j\rangle\rangle\rangle},$ (13)
where $\delta_{\langle i,j\rangle}$ is the Kronecker delta-function for nearest neighbors (i.e. it is one if $i$ and $j$ are nearest neighbors and zero otherwise), $\delta_{\langle\langle i,j\rangle\rangle}=1$ for next-nearest-neighbors, and $\delta_{\langle\langle\langle i,j\rangle\rangle\rangle}=1$ for 3rd-nearest-neighbors. The superconducting order parameters $\Delta_{ij}$ are only nonzero for nearest-neighbor sites and are either taken from solutions of the GL equations or from snapshots of the TDGL equations, using Eq. (6). To mitigate finite-size effects, results in this work are shown for a $300\times 300$ lattice. This precludes standard matrix diagonalization approaches to obtain the density of states. Rather, we use Haydock’s recursion algorithm to find the density of states at selected individual sites [62]. For a single configuration $\Delta_{ij}$, the recursion algorithm evaluates the local density of states at site $i$,
@@
where $[\ldots]^{-1}$ denotes a matrix inverse and $\delta=0.02$ is a Lorentzian broadening parameter.
@@
where the sum is over impurities, $f_{is}=\exp(-r_{is}/\kappa)/r_{is}$ for impurity $s$, $r_{is}$ is the distance between the impurity and site $i$, $\kappa$ is a characteristic length scale over which pairing is enhanced, and $\delta T^{\ast}$ sets the scale for the local enhancement. We set $\delta T^{\ast}=10T_{0}$ and $\kappa=2.8$ lattice constants, take the dopant atoms to sit a distance $0.8$ lattice constants above the plane, and set the dopant concentration to be $2.5\%$.
A.3 BKT scaling
Figure 4: Pair-phase correlation function $C_{\theta}(x,0)$ at different temperatures. Symbols show numerical results for Eq. (18) taken from TDGL data for cuts along the $x$-axis, while solid curves are fits to the asymptotic forms in Eq. (19). Fits are to the power-law form for $T\leq 0.045$ and the exponential form for $T\geq 0.55$. The data at $T=0.050$ cannot be fitted by either form.
@@ A.3 BKT scaling
[…]
@@
with $a$, $\eta$, $a^{\prime}$, and $\xi$ taken as fitting parameters. These fits are shown as solid lines in Fig. 4. Convincing fits are obtained for the power law when $T\leq 0.045$ and for the exponential when $T\geq 0.055$; however the data at $T=0.050$ is not well-reproduced by either asymptotic form, indicating that the superconducting transition temperature is $T_{c}\sim 0.050$.

### v2

@@ IV Discussion
As mentioned above, we attribute the subgap to proximity coupling between the large-gap region where the spectrum is measured, and neighboring regions with smaller gaps (see also [Sulangi:2025]). Two-gap structures like those obtained here have been observed repeatedly at low temperatures in Pb- and La-doped Bi-2201 [Boyer:2007STM, He:2014STM, Li:2021CDW, Ye:2024Emergent] and Bi-2212 [Alldredge:2008evolution, Lee:2009STM, Pushp:2009STM]. In those experiments, the subgap was interpreted as a superconducting gap and the larger gap was attributed to a non-superconducting pseudogap. However, the STM data shows that there are small-gap regions whose coherence peaks coincide with the shoulders observed in large-gap regions [Ye:2024Emergent], which is a key requirement for proximity-induced subgaps. Our calculations predict that these subgaps will fill in when thermal fluctuations wipe out the superconductivity in small-gap regions; that is, the same proximity physics that induces subgaps at low $T$ also wipes them out at high $T$. Conversely, the high-energy coherence peaks, that have been taken for “pseudogaps”, should remain above $T_{c}$, reflecting a robust and phase-coherent local superconducting patch. Unfortunately, detailed measurements of the $T$-dependent STM spectra binned according to gap size have not been reported for the overdoped materials.
@@
We remark that two-gap structures are also ubiquitous in underdoped Bi-2212 [Schmidt:2011electronic, Pushp:2009STM] and Bi-2201 [He:2014STM, Tromp:2023Puddle]. Superficially, the measured STS spectra look very much like the results presented here. However, these are complicated by the presence of a pseudogap, charge order, and possibly pair density waves [Hamidian2016]. In particular, it is likely that short-range nematic [Lee:2016nematicity], charge- [Atkinson:2018NMR] or pair-density [Choubey:2020atomic, Wang:2021scattering] order contribute to the LDOS. However, the mechanism is not understood and simple models of intertwined order generally do not reproduce the measured two-gap spectrum [Lee:2016nematicity] (see however  [Choubey:2020atomic, Wang:2021scattering]); we suggest this points to the importance of pairing inhomogeneity in underdoped cuprates. Nonsuperconducting order disappears rapidly with overdoping Bi-2201 [Li:2021CDW] and Bi-2212 [Fujita:2014simultaneous, Mukhopadhyay:2019evidence], where our model applies without modification.
We have compared in detail with Bi-2212 and Bi-2201, since high-quality STM and ARPES data is readily available for these systems, and because the distribution of local pairing interactions needed to reproduce gap maps was known from previous theoretical work [Nunner:2005dopant, melikyan2006gap] However, we expect our conclusions continue to hold qualitatively for other disordered cuprates, such as HgBa2CuO4+δ, LSCO, and Tl2Ba2CuO6+δ, since the main spectral features identified here are robust against changes in model parameters (see Fig. 5). On the other hand, stoichiometric YBa2Cu3O7, which is slightly overdoped, is both homogeneous and more three-dimensional, and is not described by our model.
@@
One of the remarkable implications of this work is that the pairing interaction is both local and varies by an order of magnitude over nm length scales. This can occur, for example, if the pairing bosons are spin fluctuations and the system is near a spin-density wave transition; in that case, inhomogeneity can create large local differences in the dynamical spin susceptibility [Romer:2012].
@@
It was recently shown that a spatially random interaction can produce a “strange-metal” linear-in-$T$ resistivity  [Bashan:2024, patel2024localization, Bashan2025:Extended] like that measured in optimally and overdoped cuprates. Our work suggests that this same strange-metal physics is responsible for anomalous features in the superconducting spectrum.
@@
To the extent that our model describes them, overdoped cuprates should exhibit aspects of both BKT and granular superconductivity. The superconducting transition in our model is essentially BKT-like, but with an inhomogeneous distribution of local phase-disordering temperatures[Benfatto:2009] and gaps[Andersen06] that broadens the transition. Our use of the Nunner model [Nunner:2005dopant] for pair inhomogeneity leads to a dilute concentration of nanoscale regions with high mean-field transition temperatures. (The importance of pair-breaking disorder for the transition in overdoped cuprates is debated  [Andersen06, Bozovic:2016Dependence, Mahmood:2019microwave, Pal:2023, Lee-Hone:2017disorder, Juskus:2024insensitivity].) In layered materials, the BKT transition is further broadened by interlayer coupling [Benfatto:2007kosterlitz].
@@
Thus, while superconductivity disappears via a smeared BKT transition, a weak diamagnetic signal will persist well above $T_{c}$ due to these nanoscale regions. This provides a natural resolution to an apparent dichotomy between magnetoresistance [Rourke:2011phase] and magnetic susceptibility [Sonier:2008, Li:2021Granular] measurements. The magnetoresistance measurements, which probe the entire sample, will depend on diamagnetic fluctuations in typical superconducting regions, while the onset temperature for a diamagnetic response in susceptibility measurements will depend on the GL temperature of the large-gap regions. We therefore believe that the anomalous spectral properties of the overdoped phase, as measured on inhomogeneous families of cuprates, can be explained by our analysis.
V Acknowledgments
@@
M.A.S. is grateful to the late J. Zaanen for numerous inspiring discussions and collaborations on related projects, and to UFIT Research Computing for providing computational resources and support that have contributed to the research results reported in this publication. W.A.A. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). This work was made possible by the facilities of the Shared Hierarchical Academic Research Computing Network (www.sharcnet.ca) and the Digital Research Alliance of Canada. P.J.H. was supported by NSF-DMR-2231821. A.K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT.
Appendix A Bogoliubov-de Gennes Calculations
A.1 Pairing Inhomogeneity
@@
The pairing interaction is assumed to be inhomogeneous because of local modulations due to out-of-plane dopant atoms [Nunner:2005dopant]. The pairing interaction $V_{ij}$ along each nearest-neighbor bond is obtained from the ansatz $V_{ij}=\frac{1}{2}(V_{i}+V_{j})$, where
$V_{i}=V_{0}+\sum^{N_{i}}_{j=1}\pm V_{I}\frac{e^{-\frac{\sqrt{|\mathbf{r}_{i}-\mathbf{r}_{j}|^{2}+r^{2}_{z}}}{\lambda}}}{\sqrt{|\mathbf{r}_{i}-\mathbf{r}_{j}|^{2}+r^{2}_{z}}},$ (3)
@@
is defined for each lattice site, $N_{i}$ is the number of impurities, the sign ($\pm$) is selected randomly for each impurity, $V_{0}=0.8$, $V_{I}=3.125$, $r_{z}=0.7$, and $\lambda=0.7$. For these calculations, we take $N_{i}=0.075L^{2}$. The created impurity realizations for these parameters are available in the Supplementary Information and are used for impurity averaging.
A.2 Density of States
@@
where we sum over $N_{k}=15\times 15$ supercells to minimize finite-size effects [Pal:2023]. In Eq. (4), the $\delta$-function is a Lorentzian of width $\eta=0.005$.
Parameter Value
$a_{0}$ 3
$T_{0}$ 0.01
$\Delta_{0}$ 0.04
$d$ 3
$\Gamma dt$ $10^{-3}$
Table 1: TDGL Parameters. All parameters are written in units of the nearest-neighbour hopping $t$, which in cuprates is $\sim 150$ meV. These parameters correspond to a filling $n=0.85$ and a low-$T$ correlation length $\xi$ of approximately 3 unit cells.
Appendix B Time-Dependent Landau-Ginzburg Calculations
B.1 Order parameter simulations
@@
[…]
@@
In this expression, $a_{0}$, $d$, $\Delta_{0}$, and $\xi_{d}$ are GL parameters, $T_{0}$ is a characteristic temperature scale for pairing, and $T_{i\alpha}^{\ast}$ is the local GL temperature below which the singlet pairing interaction along the bond $(i,i+\alpha)$ becomes attractive. Note, however, that $T_{i\alpha}^{\ast}$ is not the local critical temperature, even at the mean-field level, because the order parameter on the bond $(i,i+\alpha)$ depends on the pairing interaction in the neighborhood of the bond. The parameter $d$ determines whether the dominant pairing channel has $d$-wave ($d>0$) or extended $s$-wave ($d<0$) symmetry. In the former case, the lowest-energy GL solution for a homogeneous pairing interaction ($T_{i\alpha}^{\ast}=T_{0}$) has $d$-wave symmetry with $\eta_{x}=-\eta_{y}=\Delta_{0}$ at $T=0$. An extended $s$-wave solution is also possible, with $\eta_{x}=\eta_{y}=\Delta_{0}\sqrt{(a_{0}-2d)/a_{0}}$, however this has higher energy. Values for the GL parameters are shown in Table 1. Both the temperature scale $T_{0}$ and the pairing scale $\Delta_{0}$ are somewhat large for a quantitative description of overdoped cuprates; however, these choices help mitigate finite size effects in our simulations.
@@ B.2 Density of States
In Eq. (11), the hopping matrix elements $t_{ij}$ are the same as for the BdG calculations. The superconducting order parameters $\Delta_{ij}$ are only nonzero for nearest-neighbor sites and are either taken from solutions of the GL equations or from snapshots of the TDGL equations, using Eq. (5). To mitigate finite-size effects, results in this work are shown for a $300\times 300$ lattice. This precludes standard matrix diagonalization approaches to obtain the density of states. Rather, we use Haydock’s recursion algorithm to find the density of states at selected individual sites [Haydock:1994]. For a single configuration $\Delta_{ij}$, the recursion algorithm evaluates the local density of states at site $i$,
@@
where $\ldots]^{-1}$ denotes a matrix inverse and $\delta=0.02$ is a Lorentzian broadening parameter.
@@
where the sum is over impurities, $f_{is}=\exp(-r_{is}/\kappa)/r_{is}$ for impurity $s$, $r_{is}$ is the distance between the impurity and site $i$, $\kappa$ is a characteristic length scale over which pairing is enhanced, and $\delta T^{\ast}$ sets the scale for the local enhancement. We set $\delta T^{\ast}=10T_{0}$ and $\kappa=2.8$ lattice constants, take the dopant atoms to sit a distance $0.8$ lattice constants above the plane, and set the dopant concentration to be $2.5\%$. The resulting distribution of $T_{i\alpha}^{*}$ used in the simulations is available in the supplementary information.
B.3 BKT scaling
Figure 4: (a) Pair-phase correlation function $C_{\theta}({\bf r})$ along ${\bf r}=(x,0)$ at different temperatures. Symbols show numerical results for Eq. (16) taken from TDGL data for cuts along the $x$-axis, while solid curves are fits to the asymptotic forms in Eq. (17). Fits are to the power-law form for $T\leq 0.045$ and the exponential form for $T\geq 0.55$. The data at $T=0.050$ cannot be fitted by either form. (b) Temperature dependence of the spatially averaged phase correlation function. The average is performed over sites separated by $50<|{\bf r}|<150$ lattice constants.
@@ B.3 BKT scaling
where $\langle\ldots\rangle$ denotes a time-average. Figure 4 shows $C_{\theta}({\bf r})$ for ${\bf r}=(x,0)$ (i.e. for a cut along the $x$-axis). The correlation function is fitted to expressions for the homogeneous system, namely
@@
with $a$, $\eta$, $a^{\prime}$, and $\xi$ taken as fitting parameters. These fits are shown as solid lines in Fig. 4(a). Convincing fits are obtained for the power law when $T\leq 0.045$ and for the exponential when $T\geq 0.055$; however the data at $T=0.050$ is not well-reproduced by either asymptotic form, indicating that the superconducting transition temperature is $T_{c}\sim 0.050$.
As further confirmation, we plot the temperature dependence of $C_{\theta}({\bf r})$, averaged over separations $50<|{\bf r}|<150$ in Fig. 4(b). This figure shows that phase stiffness is lost near $T=0.055$.
Appendix C Sensitivity to Parameters
Figure 5: Density of states at $T=0.045$ for (a), (b) the $tt^{\prime}$ model shown in the main text, and (c), (d) a $t^{\prime}-t′′t^{\prime\prime}$ model. Results are shown for (a), (c) static GL calculations and (b), (d) thermally fluctuating TDGL calculations. The calculations use the same inhomogeneous GL parameters, $T_{i\alpha}^{\ast}$, as in the main text but different band structures for the calculated LDOS. The LDOS is binned such that the same sites are allocated to each bin for the two models.
Figure 6: Bin-averaged LDOS (bottom row) with inhomogeneity parameters $V_{0}=0.8$, $V_{I}=3.125$, $r_{z}=1.4$, and $\lambda=1.4$. 50 random pairing realizations and $15\times 15$ supercells are used. As with the inhomogeneity model shown in the main text, the sites are binned by the value of $\Delta_{SG}$ at $T=0.001$, and the LDOS within each bin is averaged over. The same sites are in each bin at all temperatures. Ten equal-width bins spanning $0.054<\Delta_{SG}<0.605$ are used; the legend shows the fraction of sites in each bin and darker colors correspond to lower-energy bins.
Figure 7: The binned spectral gap (a) and order parameter (b) as functions of temperature for the inhomogeneity model with $N_{i}=0.075L^{2}$, $V_{0}=0.8$, $V_{I}=3.125$, $r_{z}=1.4$, and $\lambda=1.4$. At the lowest measured temperature, each site is sorted according to its local value of $\Delta_{SG}$ and put into one of ten bins; the values of $\Delta_{SG}$ and $\Delta_{OP}$ are then averaged over each bin at each temperature.
C.1 Band-Structure Parameters
Figure 5 compares the binned LDOS for the band structure used in the main text (the $tt^{\prime}$ model) and another common tight-binding model for the cuprates (the $t^{\prime}-t′′t^{\prime\prime}$ model). For the latter model, we take t=1t=-1, $t^{\prime}=0.4$, and $t^{\prime\prime}=-0.05$. The main difference between the two models is that the van Hove singularity lies further from the Fermi level in the $t^{\prime}-t′′t^{\prime\prime}$ model, although their electron densities are the same ($n=0.85$ electrons per unit cell). Since the primary effect of doping in our model is to move the van Hove singularity towards or away from the Fermi level, these two models also provide insight into the effects of doping.
Although there are quantitative differences between the two GL calculations, Fig. 5(a) and (c), the important qualitative structures are the same: both exhibit a two-gap structure with a homogeneous subgap and an inhomogeneous spectral gap. Similarly, the TDGL calculations show identical trends, namely that the subgap fills relative to the GL gap but that the high-energy spectral gap is largely unaffected by thermal fluctuations.
C.2 Pairing Interaction Length Scale
Figure 6 shows the binned LDOS for a modified version of the inhomogeneity model used in the main text with a larger characteristic length scale. Here, we set $0.075L^{2}$, $V_{0}=0.8$, $V_{I}=3.125$, $r_{z}=1.4$, and $\lambda=1.4$ in Eq. 3. Notice that only $r_{z}$ and $\lambda$ are changed from the values used in the main text (both are doubled in this case); all other parameters are kept the same. It can be seen that much of the same striking behavior seen in the main text (i.e., the two-gap structure consisting of low-energy homogeneity and shoulders and an inhomogeneous spectral gap, and the overall anticorrelation between the spectral gap and the coherence peak) remains present even when the inhomogeneity length scale is doubled.
We show in Figure 7 the $T$-dependence of $\Delta_{SG}$ and $\Delta_{OP}$, both of which have been binned using the same procedure as in the main text. Much of what is observed in the main text is also seen here: at high temperatures, $\Delta_{SG}$ is homogeneous, whereas $\Delta_{OP}$ continues to show large variations.
Some differences between this model and the one used in the main text are: a) the spectral-gap-coherence-peak height anticorrelation is less strong in the long-ranged inhomogeneity model compared to the short-ranged one; and b) the gap persists to a larger temperature in the long-ranged model compared to the short-ranged one. However, the important qualitative features we have highlighted in the main text remain robust to changes in the inhomogeneity length scale. We defer an in-depth discussion of the length-scale dependence of the spectra and the superfluid density to a future manuscript.